OpenStudy (anonymous):

Find all solution in the interval (o, 2pi]

OpenStudy (anonymous):

\[\sin^2 x- \cos^2 x=0\]

hartnn (hartnn):

there are several ways, one of them is to use \(\sin^2x+\cos^2x=1\) so, plug in \(1-\sin^2x\) in place of cos^2 x in your equation. what u get ?

OpenStudy (anonymous):

\[2 \sin^2 x=2\]?

OpenStudy (anonymous):

do you recognize \(\sin^2(x)-\cos^2(x)\) as something else?

hartnn (hartnn):

no, where does 2 come from ? did you mean 2 sin^2 x = 1 ?

OpenStudy (anonymous):

Yes, 2 sin^2 x = 1

hartnn (hartnn):

if you know the double angle formula for cos (cos 2x =....?) then this would get easier.

hartnn (hartnn):

ok, lets continue with that, isolate sin x then, sin x =... ?

OpenStudy (anonymous):

\[\sin x= \sqrt{1/2}\] ?

hartnn (hartnn):

thats a positive root, what about negative root. \(a^2=b \implies a=\pm \sqrt b\)

OpenStudy (anonymous):

\[-\sqrt{1/2}\] then?

hartnn (hartnn):

they both \(\sin x = 1/\sqrt2 \quad \sin x=-1/\sqrt2\) do you know sin of what angle equals those ?

OpenStudy (anonymous):

\[\pi/4 & 5\pi/4 \]

OpenStudy (anonymous):

pi/4 & 5pi/4

hartnn (hartnn):

correct!

OpenStudy (anonymous):

Thanks you!!! :D

hartnn (hartnn):

welcome ^_^